If you mixed 35.0 mLmL  of 0.100 M Tris-HCl with 65.0 mLmL  of 0.200 M Tris-base, what would be the resulting pH?

pKa of Tris = 8.12
Answer #1

After mixing the solution

[tris-HCl] = 0.100M/(100ml/35ml) = 0.035M

[ tris -base ] = 0.200M / ( 100ml/65ml) = 0.130M

Henderson - Hasselbalch equation which give pH of a buffer solution is

pH = pKa + log([A-]/[HA])

[A-] = base , tris - base

[HA] = acid , tris -HCl

pKa of tris - HCl = 8.12

pH = 8.12 + log( 0.130M/0.035M)

pH = 8.12 + 0.57

pH = 8.69

Therefore,

The resulting pH = 8.69